5/2x+2/3=5/4
x= 7/30
How do I multiply the check?
5/2(7/30) + 2/3=5/4

Answer by Alan3354(69443)

5/2x+2/3=5/4
x= 7/30
How do I multiply the check?
5/2(7/30) + 2/3=5/4
-----------
35/60 + 2/3 = 5/4
7/12 + 8/12 = 5/4
15/12 = 5/4
5/4 = 5/4
===============
Or multiply thru by 60 to eliminate fractions.
35/60 + 2/3 = 5/4
35 + 40 = 75